# Stochastic Deep Learning Surrogate Models for Uncertainty Propagation in Microstructure-Properties of Ceramic Aerogels

### v1

@@ 3.2 Bayesian convolutional neural network surrogate model
where minimizing the KL divergence is equivalent to maximizing $L[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]$ over $q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ and ELBO provides a lower bound on $\ln\pi_{\text{evid}}(\boldsymbol{D})$. Considering the additive noise model, assuming Gaussian-distributed total error resulting from data noise and model inadequacy, the log-likelihood function is given by:
$\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})$ $=$ $-\frac{N_{D}}{2}\ln(2\pi\sigma_{noise})-\frac{1}{2\sigma_{noise}}\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}.$ (19)
where $R(\boldsymbol{w}_{CNN})=\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}$ is called the data misfit and $\sigma_{noise}$ is the noise variance, a hyperparameter that balances the influence of the prior (regularization term) and likelihood (data misfit term) in computing the posterior distribution. The expectation of the log-likelihood is evaluated using Monte Carlo sampling,
@@
where $\hat{\boldsymbol{w}}_{CNN}^{(j)}\sim q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ are samples drawn from the variational distribution, and $M_{mc}$ is the number of samples. In practice, the variational distribution is expressed as a product of independent Gaussian distributions following the mean-field approximation,
@@
Considering Gaussian prior, the first term in (22) has a closed-form solution, while the second term can be evaluated using (20). For numerical optimization, the ELBO from (17) is reformulated as,
@@
where $\beta=2\sigma_{\text{noise}}$ and the right-hand side of the (23) is employed in the optimization process to determine the variational parameters. To enhance the efficiency of gradient-based optimization, we applied the local reparameterization trick to convolutional layers. This approach preserves computational performance while enabling efficient updates of variational parameters. The optimization technique, commonly referred to as “Bayes by Backprop” [17], facilitates the training of BayesCNN by leveraging this reparameterization framework.
@@ 4 Numerical results
This section presents computational results from the integrated modeling framework (Figure 1), focusing on the development and evaluation of generative and BayesCNN surrogate models. Their predictive accuracy is assessed against high-fidelity methods, including the Free-surface Lattice Boltzmann Method (FSLBM) and stochastic Finite Element (FE) simulations. The linked surrogates are then used to propagate uncertainty through the synthesis-microstructure-property relationships, analyzing the impact of microstructural randomness and surrogate model errors on the reliability of the predicted quantities of interest (QoI).
@@ 4.1 Generative surrogate model of microstructure
We detail the construction of the WGAN-GP surrogate model and quantitatively evaluate its performance by comparing microstructure images generated by the WGAN-GP to the synthetic images from FSLBM simulations.
@@
Figure 4 illustrates a comparison between synthetic microstructure images of aerogels from FSLBM simulations and those generated by the WGAN-GP model. The accuracy of the WGAN-GP surrogate model is quantitatively assessed by comparing morphological attributes of generated microstructures with synthetic ones from FSLBM simulations, focusing on pore size distribution and the two-point correlation function. Figure 5(a) shows the Kernel Density Estimate (KDE) of pore diameters for 50 samples, revealing close agreement between synthetic and generated microstructures. Pore diameters were calculated using an image segmentation method that labels individual pores, determines their areas (based on pixel counts), and computes the equivalent diameter of a circle with the same area. Further details on this method are provided in [45].
@@
Figure 5: Morphological features of 50 samples from synthetic and generated microstructure images: (a) Probability distribution of pore diameters. (b) 1D plots of the two-point correlation function.
The two-point correlation function, $S_{2}(\boldsymbol{r})$, is a statistical measure characterizing the morphology of two-phase media [46], representing the probability that two points separated by $\lvert\boldsymbol{r}\rvert$ belong to the same phase. Here, we focus on the pore phase (pixels with a value of 255) to quantify pore distribution within a given distance $\boldsymbol{r}$. Figure 5(b) shows 1D plots of $S_{2}(\boldsymbol{r})$ for 50 synthetic and generated samples, with error bars capturing sample variability. The close agreement demonstrates the WGAN-GP’s ability to reproduce microstructural features consistent with physical simulations.
@@
Figure 6: Generated microstructural images of varying dimensions (in pixels) produced by the scalable generator network, corresponding to ceramic aerogel domain sizes of (a) 52.5 $\mu m$, (b) 70 $\mu m$, (c) 87.5 $\mu m$, (d) 105 $\mu m$, (e) 210 $\mu m$, and (f) 700 $\mu m$.
@@
Figure 7: Comparison of pore diameter probability distributions between synthetic 128 $\times$ 128 microstructures and scaled generated microstructures with dimensions: (a) 640 $\times$ 640 pixels (350 $\mu m$),  (b) 1280 $\times$ 1280 pixels (700 $\mu m$).
@@ 4.2 BayesCNN surrogate model of mechanical properties
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: High-fidelity training data to construct the BayesCNN surrogate model. Probability distribution of strain energy $Q^{FE}$ computed using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ and samples of three microstructure categories: small pores with average pore sizes of $6\mu m$; medium pores with average pore sizes of $18\mu m$; large pores with average pore sizes of $25\mu m$.
@@
Three categories of ceramic aerogel microstructures: small pores, medium pores, and large pores, with average pore sizes of $6\mu m$, $18\mu m$, and $25\mu m$, respectively, is taken into account (Figure 8). A total of 1,188 samples (396 per category) were simulated using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ to compute the strain energy $Q^{FE}$. For all samples, uniaxial loading conditions are applied with traction of ${t}_{y}=-10^{6}N$ being imposed downward on the top surface, while the displacement at the bottom boundary of the domain are fixed. The resulting microstructure-strain energy data was randomly divided into 90% training dataset and 10% in-distribution dataset.
@@
Figure 9: Finite element results for small pores microstructures: (a) typical high-strain-energy sample (98% of cases), and (b) anomalous low-strain-energy sample, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
@@
Figure 10: Finite element results for medium pores microstructures: (a) typical high-strain-energy sample (97% of cases), and (b) anomalous low-strain-energy sample, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
@@
Figure 11: Finite element results for large pores microstructures: (a) low-strain-energy sample (25% of cases), and (b) medium-strain-energy sample (25% of cases), and (c) high-strain-energy sample (25% of cases), showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
@@
The bimodal distribution of strain energy probabilities, illustrated in Figure 8(d), reflects the distinct mechanical behaviors exhibited within and across the three categories of ceramic aerogel microstructures. Representative FE solutions for these categories are presented in Figures 911. For the small pores category, the majority of microstructural samples (approximately 98%) exhibit strain energies ranging from $1000\times 10^{-7}$ to $1100\times 10^{-7}\,\text{mJ}$ with occasional anomalies at lower strain energy levels. As shown in Figure 9, most microstructures demonstrate mechanical deformation dominated by high-strain pathways connecting the top and bottom boundaries. In some cases, these pathways are interrupted by isolated solid-phase regions, leading to deviations in mechanical response. A similar trend is observed for the medium pores category, where strain energies predominantly range from $1150\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$, with sporadic anomalies indicating increased mechanical strength (Figure 10). In contrast, the large pores category exhibits a distinct mechanical behavior characterized by a near-normal strain energy distribution ranging from $800\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$ (Figure 11). These variations highlight the strong influence of pore size and distribution on the mechanical properties of ceramic aerogels.
@@
Figure 12: Hyperparameter selection for Bayesian inference: (a) Validation loss as a function of training epochs for various learning rates used in the optimizer. (b) Comparison of $R^{2}$ values across different $\beta$ values in both in-distribution and out-of-distribution datasets.
@@
Figures 1315 show the in-distribution prediction results of the BayesCNN model for the small pores, medium pores, and large pores microstructure classes. Each figure includes the probability distribution of high-fidelity strain energy data for the corresponding class and the three representative microstructure samples including the true strain energy value from the FE model, $Q^{FE}=\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ and the BayesCNN-predicted probability distribution, $Q^{CNN}=\mathcal{G}_{CNN}(\varphi,\boldsymbol{w}_{CNN})$, obtained using 40 samples of weight parameters $\boldsymbol{w}_{CNN}$. In all cases, the true values consistently fall within the BayesCNN-predicted distributions, demonstrating the model’s reliability for in-distribution predictions.
@@
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
@@
(b)                                                         (c)
Figure 16: Out-of-distribution predictions of the BayesCNN model: (a) Microstructure samples with average pore sizes of 12 $\mu m$ (between small pores and medium pores) and 22 $\mu m$ (between medium pores and large pores), along with the strain energy probability distribution of out-of-distribution dataset. (b) BayesCNN predictions for two microstructures with 12 $\mu m$ pores. (c) BayesCNN predictions for two microstructures with 2 $\mu m$ pores.
The accuracy and reliability of the BayesCNN predictions are further evaluated for both in-distribution and out-of-distribution dataset in Figure 17. In this figure, the prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is represented by the variance of the BayesCNN predictions, $\mathbb{V}[Q^{CNN}]$. These results indicate that the BayesCNN model successfully identifies out-of-distribution data, as reflected by lower accuracy and reliability in its predictions for such cases. The model performs well in interpolating between small pores and medium pores microstructures. However, its accuracy diminishes for out-of-distribution data with an average pore size of 22 $\mu m$, representing microstructures between medium pores and large pores.
Figure 17: The BayesCNN model’s prediction accuracy and reliability for in-distribution and out-of-distribution datasets using 10 samples. The prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is defined as $\mathbb{V}[Q^{CNN}]$.

### v3

@@ 3.2 Bayesian convolutional neural network surrogate model
where minimizing the KL divergence is equivalent to maximizing $L[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]$ over $q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ and ELBO provides a lower bound on $\ln\pi_{\text{evid}}(\boldsymbol{D})$. Considering an additive noise model, assuming Gaussian-distributed total error, the log-likelihood function is defined as:
$\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})$ $=$ $-\frac{N_{D}}{2}\ln(2\pi\sigma_{noise})-\frac{1}{2\sigma_{noise}^{2}}\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}.$ (19)
where $R(\boldsymbol{w}_{CNN})=\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}$ is called the data misfit and $\sigma_{noise}^{2}$ is a noise variance hyperparameter that regulates the relative contribution of the prior (regularization) and the likelihood (data fit) in the posterior inference. The likelihood formulation in (19), equipped with a tunable noise parameter, allows the BayesCNN surrogate to account for both aleatoric uncertainty stemming from microstructural variability in the stochastic finite element simulations, and epistemic uncertainty due to limited training data availability.
The expectation of the log-likelihood is evaluated using Monte Carlo sampling,
@@
where $\hat{\boldsymbol{w}}_{CNN}^{(j)}\sim q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ are samples drawn from the variational distribution, and $M_{mc}$ is the number of samples. To achieve computational efficiency in the high-dimensional parameter space of CNNs, where full-rank posterior approximations are intractable, the mean-field approximation is employed. It expresses the variational distribution as a product of independent Gaussian distributions,
@@
Considering Gaussian prior, the first term in (22) has a closed-form solution, while the second term can be evaluated using (20). Beyond computational convenience, the use of a Gaussian prior in BayesCNNs is motivated by the maximum entropy principle [51], which states that among all distributions with specified mean and variance, the Gaussian has the highest entropy and thus represents the least informative prior. In this work, we adopt a zero-mean Gaussian prior with variance 0.01, reflecting an unbiased assumption that positive and negative weights are equally likely before observing data. The chosen variance expresses moderate prior confidence, providing regularization to mitigate overfitting without overly constraining the capacity to learn from data [52, 53].
For numerical optimization, the ELBO from (17) is reformulated as,
@@
where $\beta=2\sigma_{\text{noise}}^{2}$ and the right-hand side of the (23) is employed in the optimization process to determine the variational parameters. To enhance the efficiency of gradient-based optimization, we applied the local reparameterization trick to the convolutional layers. This approach preserves computational performance while enabling efficient updates of variational parameters. The optimization technique, commonly referred to as “Bayes by Backprop” [22], facilitates the training of BayesCNN by leveraging this reparameterization framework. While the mean-field approximation used in this work offers a computationally tractable solution for high-dimensional BayesCNNs, its assumption of independence among weights, as defined in (21), limits its expressiveness by neglecting posterior correlations. This simplification is particularly restrictive in convolutional architectures, where weights often exhibit structured dependencies across spatial and channel dimensions. A promising alternative is the use of matrix variate Gaussian variational posteriors [54], which explicitly model covariance structures across input and output dimensions of each layer, capturing inter-parameter correlations more accurate uncertainty estimation. The applicability of this approach, along with a comparative evaluation of its predictive accuracy and computational cost relative to the mean-field approximation, will be investigated in future work within the context of microstructure–property surrogate modeling.
@@ 4 Numerical results
This section presents computational results from the integrated modeling framework (Figure 1), focusing on the development and evaluation of WGAN-GP and BayesCNN surrogate models. Their predictive accuracy is assessed against high-fidelity methods, including the Free-surface Lattice Boltzmann Method (FSLBM) and stochastic Finite Element (FE) simulations. The surrogate models are then used to propagate uncertainty through the microstructure-property relationships, analyzing the impact of microstructural randomness and surrogate model errors on the reliability of the predicted quantities of interest (QoI). The dataset generation was performed on a single CPU, with each FSLBM simulation requiring approximately 10 minutes and each FE simulation taking around 35 minutes. Model training was conducted on GPU-enabled systems, with WGAN-GP and BayesCNN training times of approximately 8 hours and 6 hours, respectively.
@@ 4.1 Generative surrogate model of microstructure
We detail the construction of the WGAN-GP surrogate model and quantitatively evaluate its performance by comparing microstructure images and morphological features generated by the WGAN-GP to the ones from FSLBM simulations.
@@
Figure 4 compares microstructure images of aerogels generated by the WGAN-GP surrogate model with those obtained from high-fidelity FSLBM simulations. To quantitatively assess the accuracy of the WGAN-GP and ensure it does not overfit or generate spurious artifacts, morphological metrics including pore size distributions and two-point correlation functions are computed and compared between the generated and reference microstructures. Figure 5(a) shows the Kernel Density Estimate (KDE) of pore diameters for 30 samples, revealing close agreement between data and generated microstructures. Pore diameters were calculated using an image segmentation method that labels individual pores, determines their areas (based on pixel counts), and computes the equivalent diameter of a circle with the same area. Further details on this method are provided in [55].
@@
Figure 5: Quantitative evaluation of the WGAN-GP generated microstructures against microstructures obtained from FSLBM simulations. Morphological features are compared across 30 microstructure samples from each model: (a) Probability distributions of pore diameters, and (b) one-dimensional two-point correlation functions.
The two-point correlation function, $S_{2}(\boldsymbol{r})$, is a statistical measure characterizing the morphology of two-phase media [56], representing the probability that two points separated by $\lvert\boldsymbol{r}\rvert$ belong to the same phase. Here, we focus on the pore phase (pixels with a value of 255) to quantify pore distribution within a given distance $\boldsymbol{r}$. Figure 5(b) shows 1D plots of $S_{2}(\boldsymbol{r})$ for 30 synthetic and generated samples, with error bars capturing sample variability. The close agreement demonstrates the WGAN-GP’s ability to reproduce microstructural features consistent with FSLBM simulations.
@@
Figure 6: Demonstration of the scalability of the WGAN-GP generator network, trained on microstructure images with a domain size of 70 $\mu m$, in synthesizing realistic microstructures across a range of domain sizes. The generator produces images with varying pixel dimensions corresponding to ceramic aerogel domains of: (a) 52.5 $\mu m$, (b) 70 $\mu m$, (c) 87.5 $\mu m$, (d) 105 $\mu m$, (e) 210 $\mu m$, and (f) 700 $\mu m$.
@@
[…]
@@ 4.2 BayesCNN surrogate model of mechanical properties
Refer to captionRefer to captionRefer to captionRefer to caption (b) Refer to captionRefer to captionRefer to captionRefer to caption (c) Refer to captionRefer to captionRefer to captionRefer to caption (d)
Figure 8: Training data of microstructure-strain energy relationship to construct the BayesCNN surrogate model: (a) Probability distribution of strain energy values $Q^{FE}$ computed using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$. Representative microstructure samples from three microstructural classes used for training: (b) small pores with average pore sizes of $6\mu m$; (c) medium pores with average pore sizes of $18\mu m$; (d) large pores with average pore sizes of $25\mu m$.
@@
Three categories of ceramic aerogel microstructures: small pores, medium pores, and large pores, with average pore sizes of $6\mu m$, $18\mu m$, and $25\mu m$, respectively, is taken into account (Figure 8). A total of 1,188 samples (396 per category) were simulated using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ to compute the strain energy $Q^{FE}$. For all samples, uniaxial loading is applied by imposing a downward traction of ${t}_{y}=-10^{6}N$ on the top surface, while the displacements at the bottom boundary are fully constrained (see Figure 3). The resulting microstructure-strain energy data was randomly divided into 90% training dataset and 10% in-distribution dataset.
@@
Figure 9: Finite element simulation results for small pores microstructures, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$ (in mm), and components of Cauchy stress $\mathbf{T}$ (in mPa): Two examples of microstructures (a) A representative microstructure (typical of 98% of samples) with a strain energy of $1098\times 10^{-7}\,\text{mJ}$; (b) An anomalous sample exhibiting a lower strain energy of $919\times 10^{-7}\text{mJ}$, attributed to isolated solid-phase regions that disrupt continuous load-bearing pathways between the top and bottom boundaries.
@@
Figure 10: Finite element simulation results for medium pores microstructures, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$ (in mm), and components of Cauchy stress $\mathbf{T}$ (in mPa): Two examples of microstructures (a) A representative microstructure (typical of 97% of samples) with strain energy value of $1265\times 10^{-7}\text{mJ}$; (b) An anomalous sample with lower strain energy ($994\times 10^{-7}\text{mJ}$).
@@
Figure 11: Finite element simulation results for large pores microstructures, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$ (in mm), and components of Cauchy stress $\mathbf{T}$ (in mPa): Three examples of microstructures (a) A low strain-energy sample; (b) A representative typical sample with a strain energy of $1049\times 10^{-7}\text{mJ}$ (characteristic of approximately 50% of the dataset); (c) A high strain-energy sample.
@@
The bimodal distribution of strain energy probabilities, illustrated in Figure 8(a), reflects the distinct mechanical behaviors exhibited within and across the three categories of ceramic aerogel microstructures. Representative FE solutions for these categories are presented in Figures 911. For the small pores category, the majority of microstructural samples (approximately 98%) exhibit strain energies ranging from $1000\times 10^{-7}$ to $1100\times 10^{-7}\,\text{mJ}$ with occasional anomalies at lower strain energy levels. As shown in Figure 9, most microstructures demonstrate mechanical deformation dominated by high-strain pathways connecting the top and bottom boundaries. In some cases, due to unexpected microstructural patterns, these pathways are interrupted by isolated solid-phase regions, leading to deviations in mechanical response. A similar trend is observed for the medium pores category, where strain energies predominantly range from $1150\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$, with sporadic anomalies indicating increased mechanical strength (Figure 10). In contrast, the large pores category exhibits a distinct mechanical behavior characterized by a near-normal strain energy distribution ranging from $800\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$ (Figure 11). These variations highlight the strong influence of pore size and distribution on the mechanical properties of ceramic aerogels.
@@
Figure 12: Selection of hyperparameters for the BayesCNN model: (a) Validation loss as a function of training epochs for various learning rates used in the optimizer. (b) Comparison of $R^{2}$ values across different $\beta$ values in both in-distribution and out-of-distribution datasets.
@@
Figures 1315 show the in-distribution prediction results of the BayesCNN model for the small pores, medium pores, and large pores microstructure classes. Each figure includes the probability distribution of high-fidelity strain energy data for the corresponding class and the three representative microstructure samples, including the true strain energy value from the FE model, $Q^{FE}=\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ and the BayesCNN-predicted probability distribution, $Q^{CNN}=\mathcal{G}_{CNN}(\varphi,\boldsymbol{w}_{CNN})$, obtained using 40 samples of weight parameters $\boldsymbol{w}_{CNN}$. In all cases, except the anomalous microstructural patterns discussed in Figures 9 and 10, the true values consistently fall within the BayesCNN-predicted distributions, demonstrating the model’s reliability for in-distribution predictions.
@@
(a)                                                         (b)
Refer to caption   Refer to caption
(c)                                                         (d)
Figure 16: Out-of-distribution predictions of BayesCNN: (a) Probability distribution of strain energy in out-of-distribution dataset consisting of 20 aerogel microstructure samples with average pore sizes of 12 $\mu m$ (interpolating between small pores and medium pores) and 22 $\mu m$ (interpolating between medium pores and large pores). (b) Representative microstructure samples for the 12 $\mu m$ and 22 $\mu m$ pore size classes. Comparison of BayesCNN-predicted strain energy with finite element simulation results for two representative out-of-distribution microstructures (c) 12 $\mu m$ pore size. (d) 22 $\mu m$ pore size.
The accuracy and reliability of the BayesCNN predictions are further evaluated for both in-distribution and out-of-distribution dataset in Figure 17. In this figure, the prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is represented by the variance of the BayesCNN predictions, $\mathbb{V}[Q^{CNN}]$. For in-distribution data, both the prediction error and uncertainty are observed to increase with average pore size. This trend correlates with the variability in the training data distributions for each microstructure class, as illustrated in Figures 1315. Specifically, the training data for the small-pore class exhibit low variance in strain energy, leading to more accurate and confident BayesCNN predictions. In contrast, the higher range of strain energy training data in the large-pore class is reflected in increased prediction error and uncertainty, indicating that the BayesCNN effectively detects the broader variability in this regime. These results indicate that the BayesCNN model successfully detects interpolation in pore size within the out-of-distribution samples, as reflected by lower accuracy (represented by prediction error $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$) and reliability (represented by the variance $\mathbb{V}[Q^{CNN}]$) in its predictions for such cases. We acknowledge that interpolation in pore size alone does not fully characterize the complexity of microstructural variability affecting mechanical response. Incorporating more expressive topological or morphological descriptors, along with principled out-of-distribution detection frameworks, such as those described in [58], could offer a more robust assessment of surrogate model generalization.
Figure 17: Prediction accuracy and uncertainty of the BayesCNN model evaluated on both in-distribution and out-of-distribution datasets using 30 samples for each average pore size. The prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is defined as $\mathbb{V}[Q^{CNN}]$.
Finally, leveraging the WGAN-GP and BayesCNN surrogate models, we perform uncertainty propagation analyses within the microstructure–property relationship, as illustrated in Figure 18. Given the demonstrated accuracy and reliability of the BayesCNN predictions for in-distribution samples (see Figure 17), the results in this figure focus exclusively on the predicted probability distributions of strain energy $\pi(Q^{CNN})$ for each microstructure class. For each class, we report the posterior mean $\mathbb{E}[Q^{CNN}]$ and the 95% credible interval $\mathcal{CI}_{95\%}[Q^{CNN}]$ of the predicted QoI, representing the uncertainty associated with microstructural variability.
$\rm{small\;pores:}$ $\mathbb{E}[Q^{CNN}]=1046.9,$ $\mathcal{CI}_{95\%}[Q^{CNN}]=[1001.2,1089.5]$
$\rm{medium\;pores:}$ $\mathbb{E}[Q^{CNN}]=1225.7,$ $\mathcal{CI}_{95\%}[Q^{CNN}]=[1173.6,1263.6]$
$\rm{large\;pores:}$ $\mathbb{E}[Q^{CNN}]=1006.15,$ $\mathcal{CI}_{95\%}[Q^{CNN}]=[915,1142.7]$
(a)                                                         (b)
Figure 18: Uncertainty propagation from microstructure to mechanical response, shown as the predicted probability distributions of strain energy (QoI) for in-distribution samples corresponding to: (a) small-pore, (b) medium-pore, and (c) large-pore microstructure classes.